HETAV PATEL

Grade 11,

A ray of light coming from the point (1,2) is reflected at a point A on the x-axis and then passes through the point (5,3). What are the co-ordinates of point A?

A ray of light coming from the point (1,2) is reflected at a point A on the x-axis and then passes through the point (5,3). What are the co-ordinates of point A? 

Grade:11

5 Answers

siddharth gupta
9 years ago
Assume the x-axis to be a  mirror.
take the image of (1,2) in x-axis i.e(1,-2).
as we know that reflected ray when extended back passes through the image of a point on incident ray i.e reflected ray passes through (1,-2) and (5,3).
...from two point form eqn of refl.ray=5x-4y-13=0.
put y=0 ...point of incidence=(13/5,0).
with regards
SIDDHARTH GUPTA 
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Harsh
6 years ago
 
Assume the x-axis to be a  mirror.
take the image of (1,2) in x-axis i.e(1,-2).
as we know that reflected ray when extended back passes through the image of a point on incident ray i.e reflected ray passes through (1,-2) and (5,3).
...from two point form eqn of refl.ray=5x-4y-13=0.
put y=0 ...point of incidence=(13/5,0).
with regards
Harsh panwar 
IIT ROORKEE
AIR 55
Yash sharma
6 years ago
Assume the x-axis to be a mirror.take the image of (1,2) in x-axis i.e(1,-2).as we know that reflected ray when extended back passes through the image of a point on incident ray i.e reflected ray passes through (1,-2) and (5,3)....from two point form eqn of refl.ray=5x-4y-13=0.put y=0 ...point of incidence=(13/5,0) Aman bansal AIR1
ankit singh
3 years ago
A ray of light passing through the point P(1,2) reflects on the x-axis at the point A and the reflected ray passes ... ∴m1=-m2. Let coordinates of point A is (a,0). Then, 2-01-a=-3-05-a ... A ray of light coming fromthe point (1, 2) is reflected at a.